OpenStudy (anonymous):

help please 6 over 5-√3?

OpenStudy (anonymous):

\(\frac{6 \times (5+\sqrt{3}} {22} \)

OpenStudy (across):

\[\frac{5}{5-\sqrt{3}}\]Rationalizing the fraction?\[\frac{5}{5-\sqrt{3}}\cdot\frac{5+\sqrt{3}}{5+\sqrt{3}}=\frac{25+5\sqrt{3}}{22}\]

OpenStudy (across):

I put 5 instead of 6 in the numerator...

OpenStudy (anonymous):

\[(5-\sqrt{3})(5+\sqrt{3})=5^2-3=22\] because \[(a-b)(a+b)=a^2-b^2\]

OpenStudy (anonymous):

thanks